# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ Appendix A Preliminary Lemmas
This section contains preliminary lemmas on RKHS and the star discrepancy, which will be useful for the subsequent proofs in Appendix˜B.
@@ A.1 Preliminary Lemmas on RKHS
$f(\theta)=\langle f(\cdot),\kappa(\cdot,\theta)\rangle_{\mathcal{H}_{\kappa}}$ (18)
at each $\theta$ pointwise (Steinwart and Christmann, 2008, Section 4.2). The reproducing property also applies to the kernel $\kappa$ itself, where it holds that $\kappa(\theta,\theta^{\prime})=\langle\kappa(\cdot,\theta),\kappa(\cdot,\theta^{\prime})\rangle_{\mathcal{H}_{\kappa}}$ at each $\theta,\theta^{\prime}$ pointwise. We have the following useful lemma on a uniform bound of functions in $\mathcal{H}_{\kappa}$.
@@
$\sup_{\theta\in\Theta}|f(\theta)|\leq 1.$ (19)
@@
$|f(\theta)|\leq\|f\|_{\mathcal{H}_{\kappa}}\|\kappa(\cdot,\theta)\|_{\mathcal{H}_{\kappa}}=\|f\|_{\mathcal{H}_{\kappa}}\sqrt{\kappa(\theta,\theta)}\leq 1$ (20)
@@
$\sup_{\theta\in\Theta}|\partial^{\alpha}f(\theta)|\leq 1$ (21)
@@
Let $I_{\alpha}$ be a set of indices $i$ s.t. $\alpha_{i}=1$. Similarly, let $I_{\setminus\alpha}$ be a set of indices $i$ s.t. $\alpha_{i}=0$. The mixed partial derivative $\partial^{\alpha}$ can be expressed as $\partial^{\alpha}=\prod_{i\in I_{\alpha}}\partial_{i}$ using the set $I_{\alpha}$. Corollary 4.36 of Steinwart and Christmann (2008) implies that
$|\partial^{\alpha}f(\theta)|\leq\|f\|_{\mathcal{H}_{\kappa}}\sqrt{\prod_{i\in I_{\alpha}}\partial_{i}\partial_{i}^{\prime}\kappa(\theta,\theta)}.$ (22)
It follows from $\|f\|_{\mathcal{H}_{\kappa}}\leq 1$ and the factorisation $\kappa(\theta,\theta^{\prime})=\prod_{i=1}^{d}\kappa_{i}(\theta_{(i)},\theta^{\prime}_{(i)})$ that
$|\partial^{\alpha}f(\theta)|\leq\sqrt{\prod_{i\in I_{\setminus\alpha}}\kappa_{i}(\theta_{(i)},\theta_{(i)})}\sqrt{\prod_{i\in I_{\alpha}}\partial_{i}\partial_{i}^{\prime}\kappa_{i}(\theta_{(i)},\theta_{(i)})}.$ (23)
@@
$\sup_{\theta\in\Theta}|\partial^{\alpha}f(\theta)|\leq\sup_{\theta\in\Theta}\sqrt{\prod_{i\in I_{\setminus\alpha}}\kappa_{i}(\theta_{(i)},\theta_{(i)})}\sqrt{\prod_{i\in I_{\alpha}}\partial_{i}\partial_{i}^{\prime}\kappa_{i}(\theta_{(i)},\theta_{(i)})}\leq 1.$ (24)
@@ A.2 Preliminary Lemmas on Star Discrepancy
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right):=\sup_{B_{0}\in\mathcal{B}_{0}}\left|\frac{\#\{\eta_{n}\in B_{0}\}}{K}-\operatorname{Vol}(B_{0})\right|$ (25)
where $\mathcal{B}_{0}$ is a set of all subboxes contained in $0,1]^{d}$ s.t. the left-boundary at every i-th coordinate is 0, and $\{\eta_{n}\in B_{0}\}$ denotes the number of points contained in a box $B_{0}\in\mathcal{B}_{0}$. Notice that the former star discrepancy $\operatorname{D}^{*}$ introduced in the main text is a generalisation of the above definition of the latter $\operatorname{D}^{\dagger}$, where they coincide with each other when $\Theta=[0,1]^{d}$.
@@
$\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{K}\right)=\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right).$ (26)
@@
Recall that $\{\eta_{n}\in B_{0}\}$ denotes the number of points contained in a box $B_{0}\in\mathcal{B}_{0}$. The number is invariant to the transform $\tau$, meaning that $\{\eta_{n}\in B_{0}\}=\#\{\tau(\eta_{n})\in\tau(B_{0})\}$, where we denote by $\tau(C)$ the image of a subset $C$ of $0,1]^{d}$ under the map $\tau$. Thus, the star discrepancy $\operatorname{D}^{\dagger}(\{\eta_{n}\}_{n=1}^{K})$ on the right-hand side can be expressed as
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)$ $=\sup_{B_{0}\in\mathcal{B}_{0}}\left|\frac{\#\{\tau(\eta_{n})\in\tau(B_{0})\}}{K}-\frac{\operatorname{Vol}(B_{0})}{\operatorname{Vol}([0,1]^{d})}\right|.$ (27)
@@
$\frac{\operatorname{Vol}(B_{0})}{\operatorname{Vol}([0,1]^{d})}=\frac{\operatorname{Vol}(\tau(B_{0}))}{\operatorname{Vol}(\tau([0,1]^{d}))}=\frac{\operatorname{Vol}(\tau(B_{0}))}{\operatorname{Vol}(\Theta)}$ (28)
@@
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)$ $=\sup_{B_{0}\in\mathcal{B}_{0}}\left|\frac{\#\{\tau(\eta_{n})\in\tau(B_{0})\}}{K}-\frac{\operatorname{Vol}(\tau(B_{0}))}{\operatorname{Vol}(\Theta)}\right|=\sup_{B\in\mathcal{B}}\left|\frac{\#\{\theta_{n}\in B\}}{K}-\frac{\operatorname{Vol}(B)}{\operatorname{Vol}(\Theta)}\right|$ (29)
@@ Appendix B Proofs
B.1 Proof of Proposition˜1
Proof.
$(\star)$ $=\int_{\Theta}\Big{|}\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\phi\left(f(\theta)\right)\right]-\phi\left(g(\theta)\right)\Big{|}^{2}d\theta=\int_{\Theta}\Big{|}\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\phi\left(f(\theta)\right)-\phi\left(g(\theta)\right)\right]\Big{|}^{2}d\theta.$ (30)
It then follows from the Jensen’s inequality that
$(\star)$ $\leq\int_{\Theta}\bigg{(}\underbrace{\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\Big{[}\left|\phi\left(f(\theta)\right)-\phi\left(g(\theta)\right)\right|\Big{]}}_{=:(*)}\bigg{)}^{2}d\theta.$ (31)
$|\phi(a)-\phi(b)|\leq\bigg{(}\sup_{c\in[a,b]}|\phi^{\prime}(c)|\bigg{)}|b-a|\leq\max\left(|\phi^{\prime}(a)|,|\phi^{\prime}(b)|\right)|b-a|$ (32)
$(*)$ $\leq\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\Big{[}\max\left(|\phi^{\prime}(f(\theta))|,|\phi^{\prime}(g(\theta))|\right)\left|f(\theta)-g(\theta)\right|\Big{]}$ (33)
$\leq\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\Big{[}\underbrace{\max\left(\exp(f(\theta)),\exp(g(\theta))\right)}_{=:(**)}\left|f(\theta)-g(\theta)\right|\Big{]}$ (34)
$(*)\leq\left(\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[(**)^{2}\right]\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\left|f(\theta)-g(\theta)\right|^{2}\right]\right)^{\frac{1}{2}}.$ (35)
Observe the following trivial inequality
$(**)^{2}=\max\left(\exp(2f(\theta)),\exp(2g(\theta))\right)\leq\exp(2f(\theta))+\exp(2g(\theta)).$ (36)
$\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[(**)^{2}\right]$ $\leq\exp(2m_{n}(\theta)+2k_{n}(\theta,\theta))+\exp(2g(\theta))$ (37)
$\leq\underbrace{\sup_{\theta\in\Theta}\exp(m_{n}(\theta)+k_{n}(\theta,\theta))^{2}+\sup_{\theta\in\Theta}\exp(g(\theta))^{2}}_{=:C}.$ (38)
Combining the bounds (35) and (38), we have the overall bound of the original term $(\star)$:
$(\star)\leq C\int_{\Theta}\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\left|f(\theta)-g(\theta)\right|^{2}\right]d\theta=C\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\int_{\Theta}\left|f(\theta)-g(\theta)\right|^{2}d\theta\right]$ (39)
B.2 Proof of Theorem˜1
@@
$I_{M+N}(f):=\left|\int_{\Theta}f(\theta)\frac{q(\theta)}{u(\theta)}u(\theta)d\theta-\frac{1}{N+M}\sum_{n=1}^{N+M}f(\theta_{n})\frac{q(\theta_{n})}{u(\theta_{n})}\right|$ (40)
@@
$d(p,\delta_{N})\leq B_{1}\sup_{\|f\|_{\mathcal{H}_{k}}\leq 1}I_{M+N}(f)+B_{2}I_{M+N}(1)+B_{3}\frac{M}{N+M}$ (41)
@@
$d(p,\delta_{N})=\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\underbrace{\left|\frac{1}{Z}\int_{\Theta}f(\theta)\frac{q(\theta)}{u(\theta)}u(\theta)d\theta-\sum_{i=1}^{N}w_{n}^{*}f(\theta_{n}^{*})\right|}_{=:(*)}\quad\text{where}\quad w_{n}^{*}=\frac{\frac{1}{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}}{\frac{1}{N}\sum_{n=1}^{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}}.$ (42)
@@
$(*)=\left|\frac{\int_{\Theta}h_{f}(\theta)u(\theta)d\theta}{Z}-\frac{\frac{1}{N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{\frac{1}{N}\sum_{n=1}^{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}}\right|$ (43)
@@
$Z_{M+N}:=\frac{1}{N+M}\sum_{n=1}^{M+N}\frac{q(\theta_{n})}{u(\theta_{n})}\qquad\text{and}\qquad Z_{N}^{*}:=\frac{1}{N}\sum_{n=1}^{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}.$ (44)
@@
$(*)$ $\leq\underbrace{\left|\frac{\int_{\Theta}h_{f}(\theta)u(\theta)d\theta}{Z}-\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{N+M}}\right|}_{=:(*_{1})}+\underbrace{\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{N+M}}-\frac{\frac{1}{N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N}^{*}}\right|}_{=:(*_{2})}$ (45)
@@
$(*_{1})$ $\leq\underbrace{\left|\frac{\int_{\Theta}h_{f}(\theta)u(\theta)d\theta}{Z}-\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z}\right|}_{=\frac{1}{Z}I_{M+N}(f)=:(*_{11})}+\underbrace{\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z}-\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{M+N}}\right|}_{=:(*_{12})}.$ (46)
@@
$(*_{12})$ $=\left|Z-Z_{N+M}\right|\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{ZZ_{N+M}}\right|\leq\frac{1}{Z}\left|Z-Z_{N+M}\right|\underbrace{\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}|h_{f}(\theta_{n})|}{\frac{1}{N+M}\sum_{n=1}^{N+M}\frac{q(\theta_{n})}{u(\theta_{n})}}}_{=:(\star)}.$ (47)
@@
$(*_{12})$ $\leq\frac{1}{Z}\left|Z-Z_{N+M}\right|=\frac{1}{Z}\left|\int_{\Theta}\frac{q(\theta)}{u(\theta)}u(\theta)d\theta-\frac{1}{N+M}\sum_{n=1}^{M+N}\frac{q(\theta_{n})}{u(\theta_{n})}\right|=\frac{1}{Z}I_{M+N}(1).$ (48)
@@
$(*_{2})$ $\leq\underbrace{\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{N+M}}-\frac{\frac{1}{M+N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N+M}}\right|}_{=:(*_{21})}+\underbrace{\left|\frac{\frac{1}{M+N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N+M}}-\frac{\frac{1}{N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N}^{*}}\right|}_{=:(*_{22})}.$ (49)
@@
$(*_{21})$ $=\left|\frac{\sum_{n=1}^{N+M}h_{f}(\theta_{n})-\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{(M+N)Z_{M+N}}\right|=\left|\frac{\sum_{\theta\in\mathcal{S}_{N+1}}h_{f}(\theta)}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}\right|\leq\frac{\sum_{\theta\in\mathcal{S}_{N+1}}q(\theta)/u(\theta)}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}$ (50)
@@
$(*_{22})$ $=\left|\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{(M+N)Z_{N+M}}-\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{NZ_{N}^{*}}\right|$ (52)
$=\frac{\left|(M+N)Z_{N+M}-NZ_{N}^{*}\right|}{(M+N)Z_{N+M}}\left|\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{NZ_{n}}\right|$ (53)
$=\frac{\left|\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})-\sum_{n=1}^{N}q(\theta_{n}^{*})/u(\theta_{n}^{*})\right|}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}\left|\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{\sum_{n=1}^{N}q(\theta_{n}^{*})/u(\theta_{n}^{*})}\right|.$ (54)
@@
$(*_{22})$ $\leq\frac{\sum_{\theta\in S_{N+1}}q(\theta)/u(\theta)}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}\leq\frac{U}{L}\frac{M}{M+N}.$ (55)
@@
$d(p,\delta_{n})$ $\leq\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{11})+\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{12})+\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{21})+\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{22})$ (56)
$\leq\frac{1}{Z}\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}I_{M+N}(f)+\frac{1}{Z}I_{M+N}(1)+2\frac{U}{L}\frac{M}{N+M}.$ (57)
@@
$\left|\int_{\Theta}g(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{K}\sum_{n=1}^{K}g(\theta_{n})\right|\leq\operatorname{Vol}(\Theta)\operatorname{V}(g)\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{K})$ (58)
@@
$V(g):=\sum_{\alpha\neq 0}\int_{\Theta^{\alpha}}\left|\partial^{\alpha}g(\theta_{*}^{\alpha})\right|d\theta^{\alpha}$ (59)

### v2

@@ Appendix A Preliminary Lemmas
This section contains preliminary lemmas on RKHS and the star discrepancy, which will be useful for the subsequent proofs in Supplement˜B.
@@ A.1 Preliminary Lemmas on RKHS
$f(\theta)=\langle f(\cdot),\kappa(\cdot,\theta)\rangle_{\mathcal{H}_{\kappa}}$ (S1)
at each $\theta$ pointwise (Steinwart2008, Section 4.2). The reproducing property also applies to the kernel $\kappa$ itself, where it holds that $\kappa(\theta,\theta^{\prime})=\langle\kappa(\cdot,\theta),\kappa(\cdot,\theta^{\prime})\rangle_{\mathcal{H}_{\kappa}}$ at each $\theta,\theta^{\prime}$ pointwise. We have the following useful lemma on a uniform bound of functions in $\mathcal{H}_{\kappa}$.
@@
$\sup_{\theta\in\Theta}|f(\theta)|\leq 1.$ (S2)
@@
$|f(\theta)|\leq\|f\|_{\mathcal{H}_{\kappa}}\|\kappa(\cdot,\theta)\|_{\mathcal{H}_{\kappa}}=\|f\|_{\mathcal{H}_{\kappa}}\sqrt{\kappa(\theta,\theta)}\leq 1$ (S3)
@@
$\sup_{\theta\in\Theta}|\partial^{\alpha}f(\theta)|\leq 1$ (S4)
@@
Let $I_{\alpha}$ be a set of indices $i$ s.t. $\alpha_{i}=1$. Similarly, let $I_{\setminus\alpha}$ be a set of indices $i$ s.t. $\alpha_{i}=0$. The mixed partial derivative $\partial^{\alpha}$ can be expressed as $\partial^{\alpha}=\prod_{i\in I_{\alpha}}\partial_{i}$ using the set $I_{\alpha}$. Corollary 4.36 of Steinwart2008 implies that
$|\partial^{\alpha}f(\theta)|\leq\|f\|_{\mathcal{H}_{\kappa}}\sqrt{\prod_{i\in I_{\alpha}}\partial_{i}\partial_{i}^{\prime}\kappa(\theta,\theta)}.$ (S5)
It follows from $\|f\|_{\mathcal{H}_{\kappa}}\leq 1$ and the factorization $\kappa(\theta,\theta^{\prime})=\prod_{i=1}^{d}\kappa_{i}(\theta_{(i)},\theta^{\prime}_{(i)})$ that
$|\partial^{\alpha}f(\theta)|\leq\sqrt{\prod_{i\in I_{\setminus\alpha}}\kappa_{i}(\theta_{(i)},\theta_{(i)})}\sqrt{\prod_{i\in I_{\alpha}}\partial_{i}\partial_{i}^{\prime}\kappa_{i}(\theta_{(i)},\theta_{(i)})}.$ (S6)
@@
$\sup_{\theta\in\Theta}|\partial^{\alpha}f(\theta)|\leq\sup_{\theta\in\Theta}\sqrt{\prod_{i\in I_{\setminus\alpha}}\kappa_{i}(\theta_{(i)},\theta_{(i)})}\sqrt{\prod_{i\in I_{\alpha}}\partial_{i}\partial_{i}^{\prime}\kappa_{i}(\theta_{(i)},\theta_{(i)})}\leq 1.$ (S7)
@@ A.2 Preliminary Lemmas on Star Discrepancy
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right):=\sup_{B_{0}\in\mathcal{B}_{0}}\left|\frac{\#\{\eta_{n}\in B_{0}\}}{K}-\operatorname{Vol}(B_{0})\right|$ (S8)
where $\mathcal{B}_{0}$ is a set of all subboxes contained in $0,1]^{d}$ s.t. the left-boundary at every i-th coordinate is 0, and $\{\eta_{n}\in B_{0}\}$ denotes the number of points contained in a box $B_{0}\in\mathcal{B}_{0}$. Notice that the former star discrepancy $\operatorname{D}^{*}$ introduced in the main text is a generalization of the above definition of the latter $\operatorname{D}^{\dagger}$, where they coincide with each other when $\Theta=[0,1]^{d}$.
@@
$\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{K}\right)=\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right).$ (S9)
@@
Recall that $\{\eta_{n}\in B_{0}\}$ denotes the number of points contained in a box $B_{0}\in\mathcal{B}_{0}$. The number is invariant to the bijective transform $\tau$, meaning that $\{\eta_{n}\in B_{0}\}=\#\{\tau(\eta_{n})\in\tau(B_{0})\}$, where we denote by $\tau(C)$ the image of a subset $C$ of $0,1]^{d}$ under the map $\tau$. Thus, the star discrepancy $\operatorname{D}^{\dagger}(\{\eta_{n}\}_{n=1}^{K})$ on the right-hand side can be expressed as
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)$ $=\sup_{B_{0}\in\mathcal{B}_{0}}\left|\frac{\#\{\tau(\eta_{n})\in\tau(B_{0})\}}{K}-\frac{\operatorname{Vol}(B_{0})}{\operatorname{Vol}([0,1]^{d})}\right|.$ (S10)
@@
$\frac{\operatorname{Vol}(B_{0})}{\operatorname{Vol}([0,1]^{d})}=\frac{\operatorname{Vol}(\tau(B_{0}))}{\operatorname{Vol}(\tau([0,1]^{d}))}=\frac{\operatorname{Vol}(\tau(B_{0}))}{\operatorname{Vol}(\Theta)}$ (S11)
@@
$\operatorname{D}^{\dagger}\left(\{\eta_{n}\}_{n=1}^{K}\right)$ $=\sup_{B_{0}\in\mathcal{B}_{0}}\left|\frac{\#\{\tau(\eta_{n})\in\tau(B_{0})\}}{K}-\frac{\operatorname{Vol}(\tau(B_{0}))}{\operatorname{Vol}(\Theta)}\right|=\sup_{B\in\mathcal{B}}\left|\frac{\#\{\theta_{n}\in B\}}{K}-\frac{\operatorname{Vol}(B)}{\operatorname{Vol}(\Theta)}\right|$ (S12)
@@ Appendix B Proofs
B.1 Proof of Theorem˜1
@@
$I_{M+N}(f):=\left|\int_{\Theta}f(\theta)\frac{q(\theta)}{u(\theta)}u(\theta)d\theta-\frac{1}{N+M}\sum_{n=1}^{N+M}f(\theta_{n})\frac{q(\theta_{n})}{u(\theta_{n})}\right|$ (S13)
@@
$d(p,\delta_{N})\leq B_{1}\sup_{\|f\|_{\mathcal{H}_{k}}\leq 1}I_{M+N}(f)+B_{2}I_{M+N}(1)+B_{3}\frac{M}{N+M}$ (S14)
@@
$d(p,\delta_{N})=\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}\underbrace{\left|\frac{1}{Z}\int_{\Theta}f(\theta)\frac{q(\theta)}{u(\theta)}u(\theta)d\theta-\sum_{i=1}^{N}w_{n}^{*}f(\theta_{n}^{*})\right|}_{=:(*)}\quad\text{where}\quad w_{n}^{*}=\frac{\frac{1}{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}}{\frac{1}{N}\sum_{n=1}^{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}}.$ (S15)
@@
$(*)=\left|\frac{\int_{\Theta}h_{f}(\theta)u(\theta)d\theta}{Z}-\frac{\frac{1}{N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{\frac{1}{N}\sum_{n=1}^{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}}\right|$ (S16)
@@
$Z_{M+N}:=\frac{1}{N+M}\sum_{n=1}^{M+N}\frac{q(\theta_{n})}{u(\theta_{n})}\qquad\text{and}\qquad Z_{N}^{*}:=\frac{1}{N}\sum_{n=1}^{N}\frac{q(\theta_{n}^{*})}{u(\theta_{n}^{*})}.$ (S17)
@@
$(*)$ $\leq\underbrace{\left|\frac{\int_{\Theta}h_{f}(\theta)u(\theta)d\theta}{Z}-\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{N+M}}\right|}_{=:(*_{1})}+\underbrace{\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{N+M}}-\frac{\frac{1}{N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N}^{*}}\right|}_{=:(*_{2})}$ (S18)
@@
$(*_{1})$ $\leq\underbrace{\left|\frac{\int_{\Theta}h_{f}(\theta)u(\theta)d\theta}{Z}-\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z}\right|}_{=\frac{1}{Z}I_{M+N}(f)=:(*_{11})}+\underbrace{\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z}-\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{M+N}}\right|}_{=:(*_{12})}.$ (S19)
@@
$(*_{12})$ $=\left|Z-Z_{N+M}\right|\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{ZZ_{N+M}}\right|\leq\frac{1}{Z}\left|Z-Z_{N+M}\right|\underbrace{\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}|h_{f}(\theta_{n})|}{\frac{1}{N+M}\sum_{n=1}^{N+M}\frac{q(\theta_{n})}{u(\theta_{n})}}}_{=:(\star)}.$ (S20)
@@
$(*_{12})$ $\leq\frac{1}{Z}\left|Z-Z_{N+M}\right|=\frac{1}{Z}\left|\int_{\Theta}\frac{q(\theta)}{u(\theta)}u(\theta)d\theta-\frac{1}{N+M}\sum_{n=1}^{M+N}\frac{q(\theta_{n})}{u(\theta_{n})}\right|=\frac{1}{Z}I_{M+N}(1).$ (S21)
@@
$(*_{2})$ $\leq\underbrace{\left|\frac{\frac{1}{N+M}\sum_{n=1}^{N+M}h_{f}(\theta_{n})}{Z_{N+M}}-\frac{\frac{1}{M+N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N+M}}\right|}_{=:(*_{21})}+\underbrace{\left|\frac{\frac{1}{M+N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N+M}}-\frac{\frac{1}{N}\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{Z_{N}^{*}}\right|}_{=:(*_{22})}.$ (S22)
@@
$(*_{21})$ $=\left|\frac{\sum_{n=1}^{N+M}h_{f}(\theta_{n})-\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{(M+N)Z_{M+N}}\right|=\left|\frac{\sum_{\theta\in\mathcal{S}_{N+1}}h_{f}(\theta)}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}\right|\leq\frac{\sum_{\theta\in\mathcal{S}_{N+1}}q(\theta)/u(\theta)}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}$ (S23)
@@
$(*_{22})$ $=\left|\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{(M+N)Z_{N+M}}-\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{NZ_{N}^{*}}\right|$ (S25)
$=\frac{\left|(M+N)Z_{N+M}-NZ_{N}^{*}\right|}{(M+N)Z_{N+M}}\left|\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{NZ_{n}}\right|$ (S26)
$=\frac{\left|\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})-\sum_{n=1}^{N}q(\theta_{n}^{*})/u(\theta_{n}^{*})\right|}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}\left|\frac{\sum_{n=1}^{N}h_{f}(\theta_{n}^{*})}{\sum_{n=1}^{N}q(\theta_{n}^{*})/u(\theta_{n}^{*})}\right|.$ (S27)
@@
$(*_{22})$ $\leq\frac{\sum_{\theta\in S_{N+1}}q(\theta)/u(\theta)}{\sum_{n=1}^{M+N}q(\theta_{n})/u(\theta_{n})}\leq\frac{U}{L}\frac{M}{M+N}.$ (S28)
@@
$d(p,\delta_{n})$ $\leq\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{11})+\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{12})+\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{21})+\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}(*_{22})$ (S29)
$\leq\frac{1}{Z}\sup_{\|f\|_{\mathcal{H}_{\kappa}}\leq 1}I_{M+N}(f)+\frac{1}{Z}I_{M+N}(1)+2\frac{U}{L}\frac{M}{N+M}.$ (S30)
@@
$\left|\int_{\Theta}g(\theta)d\theta-\frac{\operatorname{Vol}(\Theta)}{K}\sum_{n=1}^{K}g(\theta_{n})\right|\leq\operatorname{Vol}(\Theta)\operatorname{V}(g)\operatorname{D}^{*}(\{\theta_{n}\}_{n=1}^{K})$ (S31)
@@
$V(g):=\sum_{\alpha\neq 0}\int_{\Theta^{\alpha}}\left|\partial^{\alpha}g(\theta_{*}^{\alpha})\right|d\theta^{\alpha}$ (S32)